OpenStudy (anonymous):

Event A and B are independent of each other... What does independent mean here? How would it change if it said dependent. The full question continues as: The probability of event A occurring is 3x as likely as event B occurring. The probability that neither event A nor event B will occur is 7/27. What is the probability of event A occurring?

OpenStudy (anonymous):

If two events A and B are independant then P(A intersection B)=P(A)P(B)

OpenStudy (anonymous):

Just take P(B) as x P(A)=3x. Apply the condition and solve for x.

OpenStudy (anonymous):

phi the venn diagram i think is wrong, independant does not mean that it has no element in common.

OpenStudy (anonymous):

1-P(A union B)=7/27. P(A union B)=20/27. P(A)+P(B)-P(A intersection B)=20/27. x+3x-3x^2=20/27. Solve for x.

OpenStudy (anonymous):

@Phi, 15/27 is not a possible answer. Neither of you are making very much sense. I don't understand what shank means by intersection or union

OpenStudy (phi):

yep, thanks !

Directrix (directrix):

Definition: Two events, A and B, are independent if the fact that A occurs does not affect the probability of B occurring. Some other examples of independent events are: Landing on heads after tossing a coin AND rolling a 5 on a single 6-sided die. Choosing a marble from a jar AND landing on heads after tossing a coin. Choosing a 3 from a deck of cards, replacing it, AND then choosing an ace as the second card. Rolling a 4 on a single 6-sided die, AND then rolling a 1 on a second roll of the die. http://www.mathgoodies.com/lessons/vol6/independent_events.html

OpenStudy (anonymous):

Thanks directrix. Could you give some examples of dependent then?

OpenStudy (anonymous):

how can you have a union or intersection of a probability? I thought that was in sets.

OpenStudy (anonymous):

Please help me.

OpenStudy (phi):

 shankvee definitely has the right idea here